# MOMENT: A Family of Open Time-series Foundation Models

### v2

@@ E.2 Classification
$\mathtt{MOMENT_{0}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
[…]
@@
degree : 333
cache size : 200200200
max iterations : 100000001000000010000000
@@ E.3 Anomaly Detection
$\mathtt{MOMENT_{0}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
$\mathtt{MOMENT_{LP}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
initial lr : $5e-5$
Anomaly Transformer sequence length : 512512512
number of channels : 111
k : 333
anomaly ratio : 4.004.004.00
model dimensions : 512512512
number of heads : 888
embedding layers : 333
dimension of feedforward layer : 512512512
DGHL sequence length : 512512512
number of channels : 111
hidden multiplier : 323232
max filters : 256256256
kernel multiplier : 111
sub-windows : 444
size of latent z vector : 505050
number of iteration in the Langevyn dynamics inference formula : 100100100
z step size : 0.10.10.1
noise std : 0.0010.0010.001
GPT4TS sequence length : 512512512
gpt layers : 333
patch length : 111
patch stride length : 111
@@
TimesNet sequence length : 512512512
dimension of model : 161616
dimension of feedforward layer : 161616
top k : 333
number of kernels : 666
$k$-NN k : 555
@@ E.4 Imputation
$\mathtt{MOMENT_{0}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
$\mathtt{MOMENT_{LP}}$ sequence length : 512512512
patch length : 888
patch stride length : 888
initial lr : 0.00010.00010.0001
GPT4TS sequence length : 512512512
gpt layers : 333
patch length : 111
patch stride length : 111
@@
dimension of feedforward layer : 161616
TimesNet sequence length : 512512512
dimension of model : 646464
dimension of feedforward layer : 646464
top k : 333
number of kernels : 666
@@
E.5 What is MOMENT Learning?
@@
Figure 7: What is MOMENT learning? Structure in the PCA (top) and t-SNE (bottom) visualizations of the embeddings of synthetically generated sinusoids suggest that MOMENT can capture subtle trend, scale, frequency, and auto-correlation information. $\epsilon$ denotes gaussian noise with 0 mean and 0.10.10.1 standard deviation. $c$ controls the factor of interest, i.e. the power of the trend polynomial, amplitude, and frequency of the sine waves in experiments (i), (ii) & (iii), respectively.
@@
E.6 Impact of Model Size
@@ E.8 Efficiency Analysis
Appendix F Transparency Index
@@
Appendix G Results Sources
@@
Appendix H Radar Plot
We generate a radar plot (Fig. 1) to visually compare MOMENT with GPT4TS and TimesNet. The values obtained by each method for a given task are min-max normalized with respect to the other methods for each of the 5 downstream tasks. For imputation, long- and short-horizon forecasting, we report 11-1 - the normalized MSE or sMAPE for the methods on the weather and (subset of) M4 datasets, respectively. For classification and anomaly detection, we report the average accuracy and VUS-ROC of the methods across all the datasets.

### v3

@@ F.2 Classification
$\mathtt{MOMENT_{0}}$ sequence length : $512$
patch length : $8$
patch stride length : $8$
$\mathtt{SVM}$ C : $\{0.0001,0.001,0.01,0.1,1,10,100,1000,10000\}$
@@
degree : $3$
cache size : $200$
max iterations : $10000000$
@@ F.3 Anomaly Detection
$\mathtt{MOMENT_{0}}$ sequence length : $512$
patch length : $8$
patch stride length : $8$
$\mathtt{MOMENT_{LP}}$ sequence length : $512$
patch length : $8$
patch stride length : $8$
initial lr : $55e-$ 5
Anomaly Transformer sequence length : $512$
number of channels : $1$
k : $3$
anomaly ratio : $4.00$
model dimensions : $512$
number of heads : $8$
embedding layers : $3$
dimension of feedforward layer : $512$
DGHL sequence length : $512$
number of channels : $1$
hidden multiplier : $32$
max filters : $256$
kernel multiplier : $1$
sub-windows : $4$
size of latent z vector : $50$
number of iteration in the Langevyn dynamics inference formula : $100$
z step size : $0.1$
noise std : $0.001$
GPT4TS sequence length : $512$
gpt layers : $3$
patch length : $1$
patch stride length : $1$
@@
TimesNet sequence length : $512$
dimension of model : $16$
dimension of feedforward layer : $16$
top k : $3$
number of kernels : $6$
$k$-NN k : $5$
@@ F.4 Imputation
$\mathtt{MOMENT_{0}}$ sequence length : $512$
patch length : $8$
patch stride length : $8$
$\mathtt{MOMENT_{LP}}$ sequence length : $512$
patch length : $8$
patch stride length : $8$
initial lr : $0.0001$
GPT4TS sequence length : $512$
gpt layers : $3$
patch length : $1$
patch stride length : $1$
@@
dimension of feedforward layer : $16$
TimesNet sequence length : $512$
dimension of model : $64$
dimension of feedforward layer : $64$
top k : $3$
number of kernels : $6$
@@
F.5 What is MOMENT Learning?
@@
Figure 7: What is MOMENT learning? Structure in the PCA (top) and t-SNE (bottom) visualizations of the embeddings of synthetically generated sinusoids suggest that MOMENT can capture subtle trend, scale, frequency, and auto-correlation information. $\epsilon$ denotes gaussian noise with 00 mean and $0.1$ standard deviation. $c$ controls the factor of interest, i.e. the power of the trend polynomial, amplitude, and frequency of the sine waves in experiments (i), (ii) & (iii), respectively.
@@
F.6 Impact of Model Size
@@ F.8 Efficiency Analysis
Appendix G Transparency Index
@@
Appendix H Results Sources
@@
Appendix I Radar Plot
We generate a radar plot (Fig. 1) to visually compare MOMENT with GPT4TS and TimesNet. The values obtained by each method for a given task are min-max normalized with respect to the other methods for each of the 5 downstream tasks. For imputation, long- and short-horizon forecasting, we report $1-$ the normalized MSE or sMAPE for the methods on the weather and (subset of) M4 datasets, respectively. For classification and anomaly detection, we report the average accuracy and VUS-ROC of the methods across all the datasets.